OpenStudy (phebe):

Help mehhh for a medal!!!! 23. What is the value of y in the solution set of the system of linear equations shown below? y = -x + 12 4x – 2y = 36 (1 point) A-10 B-8 C-6 D-2

OpenStudy (nurali):

y = -x + 12---(1) 4x – 2y = 36 ---(2) in equation (1) y = -x + 12 x=12-y put x in equation 2 4x – 2y = 36 4(12-y)-2y=36 48-4y-2y=36 48-6y=36 -6y=36-48 -6y=-12 y=2

OpenStudy (phebe):

thanx